sadie5:

The volume of a cube is given by and the total surface area of a cube is given by , where s is the side length of the cube. If the side length of a cube is 5 inches, the volume of the cube is cubic inches and its total surface area is square inches.

sadie5:

do you know the answer

sadie5:

Joe348

FredrickStar:

yk he isn't suppose to just give out answers

sadie5:

oh did i get him in trouble

eatahgood:

can i give out an answer?

eatahgood:

ok so a cube

FredrickStar:

no u didn't get him in trouble but the rules are u are suppose to help the person figure out the answer and no u can't give out a answer u r suppose to help him figure it out

eatahgood:

what are a cubes sides?

eatahgood:

they must all be the same to be a cube otherwise it would be something else

sadie5:

In geometry, a cube is a three-dimensional solid object bounded by six square faces, facets or sides, with three meeting at each vertex. The cube is the only regular hexahedron and is one of the five Platonic solids. It has 6 faces, 12 edges, and 8 vertices.

KazGotFetee:

Volume of a cube = side^3 Total Surface Area = side^2 * 6 if side = 5 inches then Cube Volume = 5^3 = ? cubic inches

jhonyy9:

so you need to know formula for volume of a cube and formula for the total area of a cube

KazGotFetee:

so what is 5^3=?

sadie5:

yes

KazGotFetee:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sadie5 yes \(\color{#0cbb34}{\text{End of Quote}}\) Could you find the Answer to 5^3= what?

KazGotFetee:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @KazGotFetee \(\color{#0cbb34}{\text{Originally Posted by}}\) @sadie5 yes \(\color{#0cbb34}{\text{End of Quote}}\) Could you find the Answer to 5^3= what? \(\color{#0cbb34}{\text{End of Quote}}\) its basically 25 x 5 so what is this?

sadie5:

125

KazGotFetee:

Correct ok

KazGotFetee:

Total Surface Area = 5^2 * 6 and you get your Answer!

KazGotFetee:

5^2 * 6 means 25 x 6 so what is 25 x 6 =?

sadie5:

150

KazGotFetee:

CORRECT GREAT JOB!

sadie5:

thanks

KazGotFetee:

Ofc

KazGotFetee:

medal and close this Question!